# Long Video Understanding with Learnable Retrieval in Video-Language Models

### v2

@@ IV-D Ablation Studies
ActivityNet-QA 47.81/3.27 46.75/3.27 48.10/3.31 49.65/3.34 49.76/3.34
WildQA 58.00/3.30 57.45/3.18 60.58/3.31 64.82/3.39 63.44/3.39
QaEgo4D 29.74/2.43 32.42/2.41 32.04/2.42 32.51/2.45 32.81/2.42
lifeQA 33.87/2.55 37.63/2.62 38.44/2.62 38.71/2.61 37.63/2.65
Social-IQ 2.0 57.73/3.26 63.92/3.44 60.89/3.34 63.65/3.40 62.89/3.34
Average 44.84/2.89 47.86/2.91 47.99/2.92 49.92/2.96 49.19/2.95
Influence of the Time Duration of a Chunk In considering the memory limitation and video redundancy, we sample at a low frame rate (1 fps). To allow a chunk to contain dynamic temporal information, we use 4 seconds that are expected to contain temporal dynamics as the chunk unit. We study the influence of the time duration of a chunk by comparing the use of 4 frames (4 seconds) and 8 frames (8 seconds) as a chunk at the same sample rate of 1 fps. Table IX shows the performance comparison. We can see that the performance of using 4 frames/chunk is better than that using 8 frames/chunk. Firstly, 4 frames/chunk allows identification of shorter temporal segments. Secondly, it preserves more spatial details than 8 seconds per chunk since an 8-second chunk pools features over a longer period. Both the two settings outperform the baseline method Video-ChatGPT.
TABLE IX: Influence of the time duration of a chunk, evaluated in terms of accuracy (%)/score. 8f/c denotes 8frames/chunk.
Video-ChatGPT R-VLM(8f/c) R-VLM(4f/c)
ActivityNet-QA 47.81/3.27 49.24/3.31 49.65/3.34
WildQA 58.00/3.30 62.58/3.37 64.82/3.39
QaEgo4D 29.74/2.43 32.36/2.43 32.51/2.45
lifeQA 33.87/2.55 36.83/2.63 38.71/2.61
Social-IQ 2.0 57.73/3.26 62.20/3.33 63.65/3.40
Influence of the Hyperparameter $\lambda$ We determine $\lambda$ based on our previous experiences, where we usually can obtain satisfactory results by setting different losses at the same magnitude. This is a simple way to ensure that different losses are considered in optimization. We found when we set $\lambda$ as 10, the two loss terms are at the similar magnitude. We also conducted ablation study on the choose of $\lambda$ by setting it to 1, 10, and 100. Table X shows the comparison. We can see that $\lambda=10$ provides he best performance on average, slightly better than that of $\lambda=1$.
TABLE X: Ablation study on $\lambda$.
$\lambda$=1 $\lambda$=10 $\lambda$=100
ActivityNet-QA 49.70/3.32 49.65/3.34 49.06/3.33
WildQA 64.42/3.41 64.82/3.39 63.34/3.35
QaEgo4D 34.57/2.50 32.51/2.45 32.85/2.48
lifeQA 36.49/2.56 38.71/2.61 36.93/2.67
Social_IQ 2.0 62.77/3.38 63.65/3.40 60.71/3.32
Average 49.59/3.03 49.87/3.04 48.58/3.03
Figure 2: Visualization of video QA examples from QAEgo4D. The kitchen towel related to the question does not appear in the uniformly sampled video chunks. The second chunk selected by our model contains kitchen towel. Our answer states that the towel is hunging on a hook on the wall. Video-LLaMA answers incorrectly, where the towel does not appear in the first frame of the video, and it is not be placed on the countertop in front of a cutting board. Note that we use a sampled frame in a chunk to illustrate the chunk in the first two rows.
IV-E Visualization Analysis
We visualize an example from QAEgo4D in Fig. 2. More examples from QAEgo4D and WildQA are in Fig. 5, Fig. 5 in Supplementary Material B-E, with the following information.
Besides the successful cases, we present some failed cases in Fig. 9 of Supplementary Material B-E. There are two main cases of failure. One is that the retrieval did not select the correct video chunks. The other is that the retrieval correctly identified the correct video chunks, but the answer was wrong. For the latter cases, we think more powerful vision feature extractor and LLMs would alleviate the problem.
IV-F Computational Complexity
Feature Extraction Retrieval & LLM inference Total
Time (seconds) 0.14 2.42 2.56
On a single A100, we tested 120 videos with 60 seconds each from Social-IQ 2.0 and calculated the average inference time for a video, as shown in Table XI. For a video, the first part for vision feature extraction takes an average of 0.14 seconds (in parallel for 60 frames), and the second part (mainly LLM inference) takes an average of 2.42 seconds. The total time is 2.56 seconds. The retrieval by encoding the question representation, and calculating the similarity scores between the question representation and the chunks, is very fast with almost negligible time. Note that for a longer video, the time consumption for LLM inference does not increase since the input number of visual tokens to the LLM is fixed (i.e., 68$\times$5=340) in our scheme, which is favored for long video or streaming video understanding.
WildQA QaEgo4D lifeQA Social-IQ 2.0
$Avg.~{}chunks$ 19 122 20 16
$S_{vis}$ 1292 8296 1360 1088
$S_{vis}^{\prime}$ 340 340 340 340
$Save$ 69% 95% 71% 64%
[…]
$FLOPs\_Saved=\frac{S_{vis}-S_{vis}^{\prime}}{S_{tex}+S_{vis}}.$ (5)
V Conclusion
This paper is the result of an open source research project starting from 2023.
Appendix A More Details on Visual Token Extraction of a Chunk
Appendix B More Experimental Results
B-A More Details on Evaluation Metrics
In this paper, we follow the metrics of accuracy and average score as proposed by Video-ChatGPT [1] for performance evaluation, where we use ChatGPT (gpt-3.5-turbo) to assist in judging the correctness of model predictions. ChatGPT accepts questions, groundtruth answers, and the model predictions as input. For each question-answer pair, ChatGPT gives a binary judgment of “yes” or “no” to identify whether the predicted answer is correct or not, for accuracy evaluation. Moreover, an integer score of 0-5 is also given by ChatGPT to indicate how similar the prediction is to the answer. 0 represents the lowest score and 5 represents the highest score.
B-B Human Evaluation Results
We have conducted human studies in 100 samples of four test datasets on two models, the baseline scheme Video-ChatGPT, and our scheme R-VLM. We asked three people to blindly rate correct/wrong and scores and show the average evaluation results in Table R1-1. We can see that our method R-VLM achieves higher accuracy and score than Video-ChatGPT under human evaluation. The main trends from ChatGPT evaluation are consistent with our human evaluation but the values and relative gains between two schemes are different, which also indicates ChatGPT evaluation has a certain degree of reliability but are not perfect.
TABLE XIII: The results of human evaluation and ChatGPT evaluation on 100 samples, which are randomly selected from four test datasets.
Evaluator Video-ChatGPT R-VLM (Ours)
ChatGPT 48/2.98 50/3.03
Human 42.33/2.26 46.33/2.41
B-C More Ablation Studies
@@
Comparison with Q-Former-based Designs We compare our retrieval-based design with Q-former designs under our framework. 1) We use a Q-Former to globally aggregate all the chunk tokens to be the input of the LLM, where the input to Q-Former is the $L_{i}\times(N+M)$ tokens, where $L_{i}$ denotes the number of chunks, $N+M=68$ denotes the reduced number of tokens (through spatial-temporal pooling) in a chunk. The output is 340 visual tokens. There is no retrieval in this scheme, and we dub it VLM-QFormer. Table XIV shows the comparison with our scheme. We can see that our R-VLM outperforms VLM-QFormer significantly, thanks to the retrieval-based design which can exclude the inference of irrelevant visual tokens. 2) Similar to InstructBLIP [21], we use instruction-aware Q-Former to globally aggregate all the chunk tokens to be the input of the LLM. The difference with VLM-QFormer is that the instruction is also taken as the input to Q-Former to learn instruction (question)-aware visual tokens. We dub this scheme as VLM-InstructQFormer. We can see that VLM-InstructQFormer outperforms VLM-QFormer on most datasets but is inferior to our R-VLM, demonstrating the effectiveness of our retrieval-based design for extracting question-relevant visual features.
@@
TABLE XIV: Comparison with Q-Former based two designs under our framework, evaluated in terms of accuracy (%)/score.
@@
B-D More Results on Other Datasets
We also report the results on the subset of EgoSchema in Table XV. It can be seen that our R-VLM method achieves the best performance.
TABLE XV: Performance comparison with the state-of-the-art methods for video QA on the EgoSchema subset of 500 testing pairs.
@@ B-D More Results on Other Datasets
B-E More Visualization Results
@@
We visualize more examples from the QAEgo4D and WildQA datasets in Fig. 5, Fig. 5 with the following information.
@@
For the WildQA dataset, groundtruth segments usually locate at different locations in the video and last for a period of time. Although uniform sampling sometimes hits a certain groundtruth segments, there is no guarantee. In contrast, our learned retrieval can correctly hit the segments where the groundtruth segments are located on. In this way, the LLM can better understand the video and answer some detailed questions (refer to “vegetation types” in Fig. 5).
@@
In addition to successful cases, we present some failed cases in Fig.9 of Appendix B-E. There are two main cases of failure. One is that the retrieval did not select the correct video chunks. The other is that the retrieval correctly identified the correct video chunks, but the answer was wrong. For the latter cases, we think more powerful vision feature extractor and LLMs would alleviate the problem.
Figure 4: Visualization of a video QA example from QAEgo4D. The broom is small and is on the left in the red boxed image. Our R-VLM captures exactly where the broom is, i.e., on the ground. R-VLM w/Uni. does not capture the video chunks with broom and thus does not answer accurately. The answer of Video-ChatGPT is irrelevant to the question. The answer from Video-LLaMA is redundancy and tedious, where the mentioned blue jean jacket and boat actually do not appear in the video.
Figure 5: Visualization of video QA examples from WildQA. In this video, two clips show vegetation and the remaining clips show mountains, rivers, etc. Uniform sampling mainly obtains segments such as mountains and rivers rather than segments with vegetation. Therefore, only the terrain was answered, without giving vegetation types. In contrast, our retrieved chunks contain video clips of vegetation. Thus the types of vegetation are predicted correctly: trees, bushes, forest. Video-ChatGPT gives a global description and does not answer specific vegetation types.
Figure 6: A failure case from WildQA. This is a video of a person firing art. Although our R-VLM retrieved the correct chunks, it gave the wrong answer of “cook food”. This maybe caused by the visual ambiguity of the target object and the biases of the LLM.
Figure 7: A failure case from WildQA. Groundtruth chunks correspond to the chunks where three types of animals present, namely dolphins (the first chunk), seagulls, and human. Our method only retrieved the seagull and human chunks, but missed the dolphin chunk. R-VLM provided wrong answer due to the imperfect retrieval and the unsatisfactory reasoning capability of the used LLM.
Figure 8: A failure case from QAEgo4D. Our method did not find the correct chunks. Therefore, large language model did not correctly answer the question and provided hallucinated answer.
Figure 9: Visualization of failure cases from WildQA and QAEgo4D.
B-F More Discussion on Memory Cost
We have also estimated the memory cost for training different schemes when we set the batch size as 20 and use input video sequences of 80 seconds. Our scheme with $K$=5 uses similar size of memory (45.3 GB) as VideoChat-GPT (42.2 GB). In contrast, the total memory cost without using retrieval (input all the chunks, in total 1360 visual tokens) is 75.1GB, demonstrating that our retrieval-based design (45.3GB) can significantly save memory and is efficient.
B-G Discussion on the Effectiveness of Retrieval
There are no ground-truth locations of the question relevant chunks for supervision. The effectiveness of our learned chunk selection (retrieval) benefits from four aspects that promote the question (text) and vision alignment. First, the use of pretrained CLIP vision encoder and CLIP text encoder, where their features are trained to be matched/aligned. Second, the learnable MLP block (which follows the CLIP text encoder) makes it possible for the question and vision alignment, which the original CLIP pretraining (caption and vision alignment) has not been. As shown in Table VII in our main manuscript, R-VLM w/ CLIP M. denotes the use of the final CLIP class token feature of vision and text for matching instead of learnable retrieval. We can see that our final scheme R-VLM with learnable matching outperforms R-VLM w/ CLIP M. significantly. Third, the proposed soft matching (SM) loss promotes the alignment of question and vision features. R-VLM w/o SM denotes the scheme without using SM loss, which is inferior to R-VLM. Forth, the end-to-end instruction tuning (loss from the LLM) facilitates the selection of right chunks through the gradient back-propagation. Intuitively, only when the chunks are selected correctly, does the model generate the right answers. R-VLM w/o G. denotes our scheme but without back-propagating the gradient from the LLM loss. R-VLM that allows such back-propagation to the learnable MLP block outperforms R-VLM w/o G. obvious. We can see that all these factors jointly make the retrieval-based solution feasible and efficient.
Appendix C Limitation and Future Work

### v3

@@ IV-E Ablation Studies
ActivityNet-QA 47.81/3.27 46.75/3.27/93 48.10/3.31/86 49.65/3.34/80 49.76/3.34/73
WildQA 58.00/3.30 57.45/3.18/89 60.58/3.31/79 64.82/3.39/69 63.44/3.39/59
QaEgo4D 29.74/2.43 32.42/2.41/98 32.04/2.42/97 32.51/2.45/95 32.81/2.42/93
lifeQA 33.87/2.55 37.63/2.62/90 38.44/2.62/80 38.71/2.61/71 37.63/2.65/61
Social-IQ 2.0 57.73/3.26 63.92/3.44/87 60.89/3.34/76 63.65/3.40/64 62.89/3.34/52
Average 44.84/2.89 47.86/2.91/91 47.99/2.92/84 49.92/2.96/76 49.19/2.95/68
@@
Comparison with Q-Former Designs We compare our retrieval-based design with Q-Former variants. 1) VLM-QFormer aggregates all chunk tokens ($L_{i}\times(N+M)$, $N+M=68$) into 340 visual tokens without retrieval. Table IX shows R-VLM significantly outperforms VLM-QFormer, thanks to the retrieval-based design which can exclude the inference of irrelevant visual tokens. 2) VLM-InstructQFormer adds question-aware aggregation, similar to InstructBLIP [22], improving over VLM-QFormer but still underperforming R-VLM. This highlights the effectiveness of our retrieval-based design for extracting question-relevant visual features.
@@
TABLE IX: Comparison with Q-Former based two designs under our framework, evaluated in terms of accuracy (%)/score.
@@
Average 45.66/2.91 46.64/2.94 49.87/3.03
Influence of the Time Duration of a Chunk Please see Section VII-E in Supplementary for detailed analysis.
Influence of MLP Block Designs For the MLP block, we use two-layer MLP by default. We found that using deeper layers achieves similar performance (see Supplementary VII.C).
IV-F Visualization Analysis
Section VII-F in Supplementary shows visualization.
V Conclusion
Supplementary Material
VI More Details on Visual Token Extraction
@@
VII More Experiments
VII-A More Implementation Details
Following Video-ChatGPT [1], we use LLaVA 7B [18] as our base model. We utilize the pre-trained CLIP ViT-L/14 [21] as our image encoder and extract the feature from the second-to-last layer as the $h\times w$ visual tokens of a frame. We use the fine-tuned Vicuna (7B) from LLaVA as our LLM. For the text encoder, we use the pre-trained CLIP ViT-L/14 text encoder and extract the class token feature of the penultimate layer. Each of the two fully connected layers within the MLP block is equipped with 1024 neurons. The projector has 4096 neurons.
VII-B More Details on Evaluation Metrics
In this paper, we follow the metrics of accuracy and average score as proposed by Video-ChatGPT [1] for performance evaluation, where we use ChatGPT (gpt-3.5-turbo) to assist in judging the correctness of model predictions. ChatGPT accepts questions, ground-truth answers, and the model predictions as input. For each question-answer pair, ChatGPT gives a binary judgment of “yes” or “no” to identify whether the predicted answer is correct or not, for accuracy evaluation. Moreover, an integer score of 0-5 is also given by ChatGPT to indicate how similar the prediction is to the answer. 0 represents the lowest score and 5 represents the highest score.
VII-C More Results on Other Datasets
We also report the results on the subset of EgoSchema in Table X. It can be seen that our R-VLM method achieves the best performance.
TABLE X: Performance comparison with the state-of-the-art methods for video QA on the EgoSchema subset of 500 testing pairs.
@@ VII-C More Results on Other Datasets
VII-D Computational Complexity
Feature Extraction Retrieval & LLM inference Total
Time (seconds) 0.14 2.42 2.56
[…]
ActivityNet-QA WildQA QaEgo4D lifeQA Social-IQ 2.0
Avg.chunksAvg.~chunks 29 19 122 20 16
$S_{vis}$ 1972 1292 8296 1360 1088
$S_{vis}^{\prime}$ 340 340 340 340 340
SaveSave 80% 69% 95% 71% 64%
The FLOPs for LLM inference can be roughly estimated as 2PSPS https://medium.com/@dzmitrybahdanau/the-flops-calculus-of-language-model-training-3b19c1f025e4, where $P$ denotes the number of parameters, and $S$ denotes the number of tokens. The computational complexity of LLM is proportional to the number of tokens which consists of text tokens (question and answer) and visual tokens. The LLM model size $P$ is 6.7B. On the training dataset, the average number of tokens for question and answers is 80, i.e., $S_{tex}=80$. This varies on different testing datasets. For simplicity, we assume the number is the same for all the datasets. We denote the number of visual tokens as $S_{vis}$. The total number of tokens is then $S=S_{tex}+S_{vis}$. The saved computational cost (FLOPs) for LLM inference can be computed approximately as
$Saved=\frac{S_{vis}-S_{vis}^{\prime}}{S_{tex}+S_{vis}}.$ (5)
We have also estimated the memory cost for training different schemes when we set the batch size as 20 and use input video sequences of 80 seconds. Our scheme with KK=5 uses similar size of memory (45.3 GB) as VideoChat-GPT (42.2 GB). In contrast, the total memory cost without using retrieval (input all the chunks, in total 1360 visual tokens) is 75.1GB, demonstrating that our retrieval-based design (45.3GB) can significantly save memory and is efficient.
VII-E More Ablation Studies
Influence of the Hyperparameter $\lambda$ We determine $\lambda$ based on our previous experiences, where we usually can obtain satisfactory results by setting different losses at the same magnitude. This is a simple way to ensure that different losses are considered in optimization. We found when we set $\lambda$ as 10, the two loss terms are at the similar magnitude. We also conducted ablation study on the choose of $\lambda$ by setting it to 1, 10, and 100. Table XIII shows the comparison. We can see that $\lambda=10$ provides he best performance on average, slightly better than that of $\lambda=1$.
TABLE XIII: Ablation study on $\lambda$.
$\lambda$=1 $\lambda$=10 $\lambda$=100
ActivityNet-QA 49.70/3.32 49.65/3.34 49.06/3.33
WildQA 64.42/3.41 64.82/3.39 63.34/3.35
QaEgo4D 34.57/2.50 32.51/2.45 32.85/2.48
lifeQA 36.49/2.56 38.71/2.61 36.93/2.67
Social-IQ 2.0 62.77/3.38 63.65/3.40 60.71/3.32
Average 49.59/3.03 49.87/3.04 48.58/3.03
TABLE XIV: Influence of the time duration of a chunk, evaluated in terms of accuracy (%)/score. 8f/c denotes 8frames/chunk.
Video-ChatGPT R-VLM(8f/c) R-VLM(4f/c)
ActivityNet-QA 47.81/3.27 49.24/3.31 49.65/3.34
WildQA 58.00/3.30 62.58/3.37 64.82/3.39
QaEgo4D 29.74/2.43 32.36/2.43 32.51/2.45
lifeQA 33.87/2.55 36.83/2.63 38.71/2.61
Social-IQ 2.0 57.73/3.26 62.20/3.33 63.65/3.40
Average 45.43/2.96 48.64/3.01 49.87/3.03
Influence of the Time Duration of a Chunk In considering the memory limitation and video redundancy, we sample at a low frame rate (1 fps). To allow a chunk to contain dynamic temporal information, we use 4 seconds that are expected to contain temporal dynamics as the chunk unit. We study the influence of the time duration of a chunk by comparing the use of 4 frames (4 seconds) and 8 frames (8 seconds) as a chunk at the same sample rate of 1 fps. Table XIV shows the performance comparison. We can see that the performance of using 4 frames/chunk is better than that using 8 frames/chunk. Firstly, 4 frames/chunk allows identification of shorter temporal segments. Secondly, it preserves more spatial details than 8 seconds per chunk since an 8-second chunk pools features over a longer period. Both the two settings outperform the baseline method Video-ChatGPT.
Influence of MLP Block Designs For the MLP block, we use two-layer MLP on text encoder by default. As shown in Table XV, we found that using deeper layers (six-layer MLP) achieves similar average performance.
As discussed in Section Section III.B, we add an MLP block after the text encoder to facilitate the matching. Another alternative design is to apply an MLP to the vision encoder for retrieval. We added an ablation comparison of adding MLP after vision encoder (R-VLM (2-layer MLP on vision)) vs. adding MLP after text encoder (R-VLM (2-layer MLP on text)). Table XV shows that similar average performance is achieved.
TABLE XV: Ablation study on MLP block designs: two layers vs. six layers after the text encoder, and MLP after the vision encoder. Performance is evaluated in terms of accuracy (%)/score.
R-VLM R-VLM R-VLM
(2-layer MLP (6-layer MLP (2-layer MLP
on text) on text) on vision)
ActivityNet-QA 49.65/3.34 49.02/3.32 49.71/3.25
WildQA 64.82/3.39 62.83/3.34 64.36/3.36
QaEgo4D 32.51/2.45 32.29/2.44 31.75/2.42
lifeQA 38.71/2.61 42.30/2.65 39.33/2.59
Social-IQ 2.0 63.65/3.40 63.31/3.39 61.82/3.36
Average 49.87/3.03 49.95/3.03 49.39/3.00
Figure 3: Visualization of video QA examples from QAEgo4D. The kitchen towel related to the question does not appear in the uniformly sampled video chunks. The second chunk selected by our model contains kitchen towel. Our answer states that the towel is hunging on a hook on the wall. Video-LLaMA answers incorrectly, where the towel does not appear in the first frame of the video, and it is not be placed on the countertop in front of a cutting board. Note that we use a sampled frame in a chunk to illustrate the chunk in the first two rows. GPT-4o correctly points out where the towel is.
Figure 4: Visualization of a video QA example from QAEgo4D. The broom is small and is on the left in the red boxed image. Our R-VLM captures exactly where the broom is, i.e., on the ground. R-VLM w/Uni. does not capture the video chunks with broom and thus does not answer accurately. The answer of Video-ChatGPT is irrelevant to the question. The answer from Video-LLaMA is redundancy and tedious, where the mentioned blue jean jacket and boat actually do not appear in the video. GPT-4o gave the correct answer, i.e., the broom is on the ground.
Figure 5: Visualization of a video QA example from QAEgo4D. We can see that the gray car does not appear in the uniformly sampled video chunks. Our R-VLM correctly answers that the car was parked in the parking lot (outdoors), but R-VLM w/Uni.’s answer was the garage (indoors). Video-LLaMA does not answer where the car is and the ground-truth frames do not appear in the frame 100. Video-ChatGPT made the similar mistake as R-VLM w/Uni. GPT-4o had an illusion that the car was parked on a gravel road, but the road was flat, so it gave an inaccurate answer.
Figure 6: Visualization of video QA examples from WildQA. In this video, two clips show vegetation and the remaining clips show mountains, rivers, etc. Uniform sampling mainly obtains segments such as mountains and rivers rather than segments with vegetation. Therefore, only the terrain was answered, without giving vegetation types. In contrast, our retrieved chunks contain video clips of vegetation. Thus the types of vegetation are predicted correctly: trees, bushes, forest. Video-ChatGPT gives a global description and does not answer specific vegetation types. GPT-4o gave correct answer.
Figure 7: Visualization of a video QA example from WildQA. The uniform sampling miss the chunks for sharpening process in ground-truth segments (at the beginning of video). As a result, LLM does not see the knife running along the leather, and only see the knife and some delicate small objects. Therefore, R-VLM w/Uni. mistakenly thought that this individual was carving patterns or making designs. Our retrieved chunks retain the process of the knife running on the leather and therefore R-VLM gives the correct answer. Both Video-LLaMA and Video-ChatGPT answered that people are cutting leather with a knife to make art, rather than sharpening the knife. GPT-4o thinks the knife is carving leather, which is inconsistent with the correct answer: sharpening the knife.
Figure 8: Visualization of a video QA example from WildQA. In this video, collecting the kindling takes a short time, while placing the tinder on the stones takes a longer time. Uniform sampling makes LLM think that there is no process of collecting kindling and output the wrong answer of “digging a hole”. Our R-VLM identified the relevant chunks of “gathering” even though those chunks only take a small duration in the entire video, generating correct answer. Video-LLaMA’s prediction is not accurate since in fact the man did not gather deadwood and leaves in the video. GPT-4o gave the correct answer on gathering kindling.
(a) A failure case from WildQA. This is a video of a person firing art. Although our R-VLM retrieved the correct chunks, it gave the wrong answer of “cook food”. This maybe caused by the visual ambiguity of the target object and the biases of the LLM.
(b) A failure case from WildQA. ground-truth chunks correspond to the chunks where three types of animals present, namely dolphins (the first chunk), seagulls, and human. Our method only retrieved the seagull and human chunks, but missed the dolphin chunk. R-VLM provided wrong answer due to the imperfect retrieval and the unsatisfactory reasoning capability of the used LLM.
(c) A failure case from QAEgo4D. Our method did not find the correct chunks. Therefore, large language model did not correctly answer the question and provided hallucinated answer.
Figure 9: Visualization of failure cases from WildQA and QAEgo4D.
Figure 10: Visualization of video QA examples from QAEgo4D. For the curves, the red pointed stars indicate the video clip selected by R-VLM (red line), and the blue dots indicate the video clip selected by R-VLM w/o SM (blue line). The green dotted line indicates the ground truth question-related clips. After adding SM Loss training, our model R-VLM can capture the corresponding video clips more accurately.
Figure 11: Visualization of video QA examples from WildQA. For the curves, the red pointed stars indicate the video clip selected by R-VLM (red line), and the blue dots indicate the video clip selected by R-VLM w/o SM (blue line). The green dotted line indicates the ground truth question-related clips. After adding SM Loss training, our model R-VLM can capture the corresponding video clips more accurately.
VII-F Visualization Analysis
@@
Overall Visualization We visualize examples from QAEgo4D in Fig. 4, Fig. 4, Fig. 5, and examples from WildQA in Fig. 6, Fig. 7, Fig. 8, with the following information.
@@
Failure Case Visualization In addition to successful cases, we present some failed cases in Fig.9. There are two main cases of failure. One is that the retrieval did not select the correct video chunks. The other is that the retrieval correctly identified the correct video chunks, but the answer was wrong. For the latter cases, we think that more powerful vision feature extractor and LLMs would alleviate the problem.
Visualization on Influence of SM Loss We added visualizations showing similarity curves of chunk-question embeddings before and after SM loss training in Fig. 11 and Fig. 11. Note that for the curves, the red pointed stars indicate the video clip selected by R-VLM (red line), and the blue dots indicate the video clip selected by R-VLM w/o SM (blue line). The green dotted lines indicate the ground truth question-related clips. We can see that, in general, after adding SM Loss training, the model can capture the corresponding video clips more accurately.
VII-G Human Evaluation Results
We have conducted human studies in 100 samples of four test datasets on two models, the baseline scheme Video-ChatGPT, and our scheme R-VLM. We asked three people to blindly rate correct/wrong and scores and show the average evaluation results in Table XVI. We can see that our method R-VLM achieves higher accuracy and score than Video-ChatGPT under human evaluation. The main trends from ChatGPT evaluation are consistent with our human evaluation, which indicates the reliableness of ChatGPT evaluation.
TABLE XVI: The results of human evaluation and ChatGPT evaluation on 100 samples, which are randomly selected from four test datasets.
Evaluator Video-ChatGPT R-VLM (Ours)
ChatGPT 48.00/2.98 50.00/3.03
Human 42.33/2.26 46.33/2.41
VIII Discussion
VIII-A Rationale for Model Selection
Here, we discuss the rationale for choosing specific models like MLP, CLIP image encoder, text encoder, and LLM.
CLIP vision and text encoders We use CLIP because it supplies robust, semantically aligned embeddings across modalities, thanks to its contrastive pretraining on 400M image-text pairs. This joint embedding space enables effective cross-modal alignment and efficient chunk scoring without expensive fine-tuning. CLIP’s encoders have been widely adopted in Video Language Modeling for VQA.
Large Language Model (LLM) For understanding retrieved chunks and generating answers, LLMs offer strong reasoning and language capabilities. We chose LLaVA as our base model, which uses the popular LLM of LLaMA, enabling the process of natural language and visual context effectively.
MLP for retrieval A learnable two-layer MLP following the text encoder provides a lightweight and efficient way to score chunk relevance for retrieving question relevant video chunks.
VIII-B About Hour-long Videos
We tested some videos from the popular long video benchmark LVBench (an average length of 68 minutes). We found that (1) For question-video pairs with long certificate length (shortest video span that contains sufficient evidence to answer the question correctly), our model can capture the corresponding video chunks and generate answer correctly. For example, given the question: “What does the girl use to find the person she is looking for in the video? (A) Mask (B) Tail (C) Phone (D) Car”. Our model predicted the correct answer “A”. (2) For the pairs whose answers only exist in very short clips (1-2 chunks), since the average number of chunks is more than 1000, it becomes hard for our model to select the correct clips, making it difficult to give a correct answer. More efforts are still needed to deal with such hour-long videos and we leave as future work.
IX Limitation and Future Work